# The Bathtub of European AI Governance: Identifying Technical Sandboxes as the Micro-Foundation of Regulatory Learning

### v1

@@ 3.2. AITS in the Bathtub
Regulatory learning in a data-driven adaptive legislative framework depends on the availability of evidential data generated at the micro level. Under the AI Act, compliance with legislative obligations is assessed through concrete technical evaluations. The data produced by these assessments therefore constitutes the primary stimulus for learning and adaptation at the meso and macro levels.
However, as shown in the previous section, the regulatory learning space defined by the AI Act involves a dense network of actors and activities. If the evidential data produced at the micro level were to emerge from heterogeneous, uncoordinated methodologies and infrastructures, its aggregation, comparison, and interpretation would be severely constrained. Rather than enabling learning, such fragmentation would increase administrative overhead and obscure regulatory signals.
To support meaningful regulatory learning, common technical frameworks, shared infrastructures, and aligned assessment practices are therefore required to structure micro-level evidence in a way that can be operationalised at the meso and macro levels.
At the policy level, the AI Act explicitly emphasises AIRS as a formal instrument for experimentation and supervision, and implicitly as a channel for regulatory learning. In (Buscemi et al., 2025), Buscemi et al. examined AIRS in detail, with a particular focus on the role and technical requirements of AITS operating within the framework established by the AI Act. While regulatory learning was acknowledged as a natural by-product of AIRS, its mechanisms, scope, and implications were not analysed in depth.
Existing literature on regulatory learning under the AI Act largely concentrates on this single pathway, framing AIRS as the primary feedback mechanism connecting national CA to the Commission (Ahern, 2025a). However, the AI Act provides for additional compliance pathways that also generate structured technical evidence and may therefore play an instrumental role in regulatory learning. Notably, Article 43 allows conformity assessment to be conducted either through internal controls or via a notified body.
On this basis, we identify three distinct scenarios through which micro-level evidence is produced and can contribute to regulatory learning: (1) self-assessment by a provider or deployer, (2) assessment within an AIRS, and (3) conformity assessment performed by a notified body.
@@
Figure 2. The flow of information in the self-assessment by an SME. While the regulatory learning here is primarily internal to the SME and constrained to the micro level, its participation in structures such as standardisation processes and the Advisory Forum allows for their information and experience to also be passed to the meso and macro levels.
The flow of this scenario, with respect to the regulatory learning process, is shown in Figure 2. The legal requirements represent the macro-level pressure exerted on the micro-level SME to guarantee compliance. For the SME, legal requirements such as fairness are abstract with respect to their AI system. Hence, this step in the process requires careful interpretation by the SME in order to translate the legal requirements into appropriate technical requirements and a subsequent selection of corresponding assessment solutions. As provided for by the AI Act, said translation should be facilitated by regulatory instruments such as the harmonized standards that enable presumption with conformity.
The current absence of such regulatory instruments gives this selection particular potential for impact on regulatory learning. At the micro level, the iterative assessment of the AI system and the accompanying selection of tools with a consistent, reproducible methodology in the AITS makes the development of the AI system evident and transparent, potentially helping with interpretation of the legal requirements and the assessment results.
At the same time, the scenario is not necessarily constrained to the micro level and the data generated may be passed onto the meso level. In fact, the SME may be a member of the Advisory Forum as provided for by Article 67(2), thus creating a direct feedback channel at the meso level to integrate the learned micro-level evidence. As the Advisory Forum collaborates with the AI Board, AI Office, and, if requested, with the Commission, the generated evidence can also reach the macro level. In addition, the SME may also participate in the standardization processes of CEN and CENELEC (62), allowing them to directly affect the design of European standards, which constitutes another feedback loop within the regulatory learning process. In these processes, the assessment tool, method, and metrics selection helps the standardisation bodies detect potential candidates for adoption in standards.
AI Regulatory Sandbox. Some providers, especially start ups and SMEs, do not have the resources or the legal expertise to perform the self-assessment and thus need external support. For those cases, and as a formal mechanism of regulatory learning, article 57 of the AI Act defines the previously mentioned AIRS.
@@
In contrast to the self-assessment, however, this scenario represents not an indirect, but a direct, legally formal flow of information from the micro to the meso level. If the AITS methodology is employed on a consistent basis in AIRS engagements by the MSA, engagements become comparable and the MSA themselves gather evidence to improve their understanding of the pipeline from high-level legislation to technical operationalisation. As the MSA are tasked with sharing the collected data in their collaboration with the AI Board and the AI Office, it proliferates through the meso level, ultimately reaching the macro level. With the number of AIRS engagements growing, the machine-readable data generated within the AITS supports aggregation and allows scalable analysis at the meso and macro level. For the AI Office, it becomes particularly valuable in the design of guidelines and Codes of Practice, whereas the Commission may use insights to judge the appropriateness of standards to receive legal force through delegated acts.
@@
Conformity assessment by a notified body. The third, and last, scenario that we consider here consists in the certification process performed by a NB, certified by a national NA. This scenario, anchored in the AI Act via articles 43(1)(b), 43(3), and annex VII, can be considered a continuation of the two previous scenarios. Since the certification process consists uniquely in a review of the quality management system implemented and the technical documentation drafted by a provider or deployer, either self-assessment or an engagement similar to the AIRS has to precede it. Nevertheless, due to their responsibilities in reporting to the MSA, the NB serve as a first meso-level aggregator passing on the collected evidence. As a first information processor, they can perform a preliminary analysis of sector-specific data, improving clarity in the later stages of the regulatory learning process. For that to happen efficiently and at scale, manual, text-based compliance reports need to be replaced by specifically designed, standardised machine-readable formats, employed by the AITS.
4. Towards Modular AITS as Regulatory Learning Tools
The previous section examined the role of AITS as a source of micro-level evidence in support of regulatory learning across different AI assessment scenarios. That discussion remained at an abstract, functional level.
@@ 4. Towards Modular AITS as Regulatory Learning Tools
This section therefore focuses on the infrastructural and formal requirements needed to operationalize AITS as a foundational component of the EU AI regulatory landscape. In light of the absence of mandated machine-readable formats in the AI Act (Schrepel, 2025), particular attention is given to solutions that enable scalable and (semi-)automated processing, with the aim of reducing the regulatory burden associated with manual documentation. While not exhaustive, the components discussed below are considered essential to achieving the flexibility required for this purpose.
@@
Extensible Formal Configuration Language. In (Buscemi et al., 2025), a Domain-Specific Language (DSL) was proposed as a formal, machine-readable specification framework for AITS. From a legal perspective, this DSL can be understood as a standardised formal template that structures how technical requirements, assessment procedures, and evidential outputs are specified and documented. Acting as the common specification framework for AITS, the DSL encodes assessment logic in a way that mirrors how AI systems are designed and evaluated in practice. It distinguishes between a stable core layer, which captures requirements that apply horizontally across sectors, and sector-specific extensions that introduce additional obligations relevant to particular application domains. This layered structure follows a piggybacking design pattern (Mernik et al., 2005), whereby sector-specific provisions are formally anchored to a shared core. As a result, domain-specific requirements remain legally and technically traceable to the horizontal foundations defined in the core DSL. This traceability supports regulatory learning both within individual sectors and across the regulatory system as a whole.
Internal Unified Data Model. A core requirement of modular AITS is the ability to integrate assessment solutions from an open catalogue, including both open-source and proprietary tools, and to present their results through harmonised reporting and visualisation mechanisms. These elements are essential for conformity assessment and regulatory learning, as they enable auditability, comparability, and ex post review of the technical evidence produced.
@@
In practice, assessment tools reflect the diversity of AI systems covered by the AI Act, ranging from traditional machine learning systems to generative models (European Commission, 2025a). Differences in system architectures, inputs, and outputs lead to heterogeneous evaluation methods and metrics, which complicates the harmonised storage and interpretation of results within the AITS. To preserve evidentiary value, the underlying data infrastructure must therefore support flexible data models and fine-grained access, including instance-level evidence where aggregate metrics would be insufficient (Burnell et al., 2023).
Because assessment tools implement distinct evaluation logics, their integration requires not only architectural flexibility but also an explicit mapping between tool-specific outputs and a common evidentiary structure. To support a scalable ecosystem of assessment solutions while preserving human oversight, this mapping should be supported through semi-automated, user-validated mechanisms, enabling a plug-in architecture consistent with human-in-the-loop requirements.
Assessment tool documentation. Existing surveys and taxonomies of AI trustworthiness assessment reveal a highly fragmented evaluation landscape (Huang et al., 2025; Campagner et al., 2023; Newman, ; Braga et al., 2025). This fragmentation raises not only technical concerns but also legal ones, as divergent methodologies, metrics and documentation practices complicate the interpretation, comparability and auditability of assessment outcomes. Assessment tools may address the same legal requirement while relying on different underlying metrics or assumptions, thereby undermining consistent compliance assessment and legal certainty.
@@
The mere integration of such tools as plug-ins, while essential for flexibility, is therefore insufficient to meet the requirements of transparency and traceability. A complementary approach lies in the adoption of standardised reporting formats based on a shared vocabulary for assessment solutions. Comparable instruments already exist in the form of model cards and data sheets (Mitchell et al., 2019; Gebru et al., 2021; Roman et al., 2024), with more recent proposals such as AI cards, use case cards and AI product cards responding to documentation needs arising under the AI Act (Golpayegani et al., 2024; Hupont et al., 2024; Celepija et al., 2025).
@@
When accompanied by a machine-readable representation, such standardised formats enable the structured cataloguing of assessment solutions and support their consistent application across regulatory contexts. This, in turn, facilitates automated or semi-automated tool selection aligned with applicable legal requirements, reducing barriers to effective conformity assessment.
At the meso and macro levels, standardised reporting also provides a basis for regulatory learning. By enabling systematic analysis of how assessment tools are applied in practice, competent authorities, the AI Office and standardisation bodies can derive evidence to inform guidance, standards and recommended practices, thereby strengthening consistency and legal certainty over time.
Shared and Standardised Metrics Vocabulary. While the aforementioned DSL would constitute a common, machine-readable framework for the AITS specification, the usefulness of its structure ultimately depends on a consistent, shared vocabulary. If the specified assessment logic within the DSL configuration does not use the same terminology across different sandbox instances, e.g.  if different definitions of the same metric are employed, consistent and significant comparison of assessments becomes impossible. Especially the definition of such metrics, specific to the type of AI system and legal requirement they address, through machine-readable ontologies in open standard formats such as the Resource Description Framework (RDF) (57) would allow mapping between assessment tools and methodologies.
@@
In literature, significant steps in this direction have been made through ontologies such as AIRO, an ontology of risks in AI systems (Golpayegani et al., 2022) and semantic mapping between trustworthiness requirements and standards (Hernandez et al., 2025). However, the proposed ontologies remain high-level, adhering to the abstract definitions of the AI Act, with a closer link to operational technical requirements still missing. In (Di Scala et al., 2025), a conceptual language bridging the normative legal requirements to the functional components of an AI system is proposed. A combination of such semantic frameworks with standardised assessment tool documentation could drive the semi-automated instantiation of AITS and the creation of auditable and machine-readable assessment reports with the needed consistency for analysis and regulatory learning at scale.
@@ 5. Discussion
The theoretical framework described in Section 3.1 establishes the micro level as the location of evidence generation, with the subsequent sections elaborating on the AITS as the critical technical infrastructure to complement formal regulatory mechanisms such as the AIRS. In this section, we discuss the implications that the AITS as a standardised assessment tool would entail for regulatory learning, namely its strengths and associated socio-technical and socio-political challenges.
@@
First, a fundamental issue lies in learning at each level taking place at different speeds, with the learning at the micro level following the fast cycles of software development and thus vastly outpacing the learning at the macro level, itself constrained by the defined legislative review periods of up to four years (Article 112) (Ahern, 2021). The meso level, tasked with bridging this gap through harmonisation, currently risks becoming the bottleneck due to additional delays in the standardisation process. In this context, the AITS acts as a technical buffer decoupling innovation speed from regulatory inertia. While the AI Act prescribes manual documentation, impeding scalable analysis, the proposed machine-readable formats of the AITS provide a basis for future automated evidence aggregation. Due to data normalisation at the source, i.e. the micro level, through its internal unified data model, the AITS streamlines the aggregation and interpretation of reported results at the meso level, alleviating the burden of interpreting manually reported documents at scale. This would in turn enable the meso level to identify technical trend with an agility that the traditional legislative cycles cannot match.
@@
A further, fundamental, challenge lies in the definition of adequate metrics and assessment methodologies to accurately capture the impact of an AI system on the fundamental rights the AI Act aims to codify. Abstract requirements such as fairness are inherently difficult to define at a technical level. This puts particular demands on the regulatory learning cycle as technical developments have to be taken into account when operationalising legislation. The AITS, through the plug-in architecture can alleviate some of this pressure by enabling consistent experimentation and validation of diverse metrics. This way, empirical discovery of potentially adequate metrics, methods, and methodologies becomes available at a large scale, not just in the conformity assessment, but in the learning processes. Especially the meso actors, tasked with the translation of the legal requirements, can benefit from said discovery in the design of the sector-specific, vertical regulatory instruments such as Codes of Practice and the harmonised standards. At the same time, this particular concern calls for caution, as scholars have already warned against standards aiming at becoming too granular through the definition of, e.g., overly specific thresholds (Gornet and Maxwell, 2024).
@@
Beyond this normative challenge, the ambition of an adaptive governance model is not merely a technocratic matter in which learning naturally takes place from data. Instead, regulatory learning is inherently political, subject to significant influence from the interests of the involved stakeholders and decision makers. At the legislative level, this manifests as a question of political willingness. This issue is further compounded at the meso level where the data is aggregated and filtered before passing it to the macro level, especially in the standardisation efforts intended to provide vertical governance. As these standardisation processes are often dominated by largy industry players with significant resources, there is a non-negligible risk of regulatory capture, where the interpretation of the legal requirements as technical operationalisations is skewed by optimising for commercial feasibility and profit instead of the robust protection of fundamental rights (Metcalf, 2025). In particular the interests of the micro-level actors, such as SMEs, should be given more weight (Kilian et al., 2025), a step that the AI Act has already taken through their inclusion in the regulatory learning process via the Advisory Forum. As the AITS is a technical infrastructure enabling transparent, reproducible, and consistent assessments, it is decidedly not prescriptive. It is a tool enabling generation of data, with political issues out of its scope. However, the AITS methodology and subsequent semi-automatic reporting would alleviate the need for resources and effectively democratise the aggregation layer through potentially higher impact by smaller stakeholders. Requiring the use of the evidential data in the standardisation process could potentially transform the decision making by giving equal weights to the input from small actors as the one from large industry players.
@@
Finally, evidence generation by providers and deployers is insufficient as the data needs to be shared to enable learning at the higher levels. This raises the crucial question concerning the incentive for industry stakeholders to allow access to their logged data. On the one hand, privacy and data protection concerns naturally arise as companies will not want to reveal internal information. On the other hand, the AITS and its generated data runs the risk of becoming self-incriminating for companies instead of supporting their compliance endeavours. To mitigate such liability concerns, early stage AITS findings could, similarly to the AIRS engagements, be legally treated as preliminary learning signals instead of grounds for potential sanctions, acknowledging the iterative nature of development and assessment. This could lead to actors at the micro level being more open towards sharing their results arising from self-assessment beyond just going through the necessary steps for conformity assessment, as it would signal that their input is explicitly desired to review the legislation and regulations. Furthermore, by reducing the resources required in drafting evidence of compliance, the AITS could see improved adoption in the industry, fuelling both enforcement and learning at the same time.

### v2

@@ 3.2. AITS in the Bathtub
Regulatory learning in a data-driven adaptive legislative framework depends on the availability of evidential data generated at the micro level. Under the AI Act, compliance with legislative obligations is assessed through concrete technical evaluations. The data produced by these assessments therefore constitutes the primary stimulus for learning and adaptation at the meso and macro levels. However, as shown in the previous section, the regulatory learning space defined by the AI Act involves a dense network of actors and activities. If the evidential data produced at the micro level were to emerge from heterogeneous, uncoordinated methodologies and infrastructures, its aggregation, comparison, and interpretation would be severely constrained. Rather than enabling learning, such fragmentation would increase administrative overhead and obscure regulatory signals.
To support meaningful regulatory learning, common technical frameworks, shared infrastructures, and aligned assessment practices are therefore required to structure micro-level evidence in a way that can be operationalised at the meso and macro levels. At the policy level, the AI Act explicitly emphasises AIRSes as a formal instrument for experimentation and supervision, and implicitly as a channel for regulatory learning. In (Buscemi et al., 2025), Buscemi et al. examined AIRSes in detail, with a particular focus on the role and technical requirements of AITSes operating within the framework established by the AI Act. While regulatory learning was acknowledged as a natural by-product of AIRSes, its mechanisms, scope, and implications were not analysed in depth.
Existing literature on regulatory learning under the AI Act largely concentrates on this single pathway, framing AIRSes as the primary feedback mechanism connecting national CA to the Commission (Ahern, 2025a). However, the AI Act provides for additional compliance pathways that also generate structured technical evidence and may therefore play an instrumental role in regulatory learning. Notably, Article 43 allows conformity assessment to be conducted either through internal controls or via a notified body. On this basis, we identify three distinct scenarios through which micro-level evidence is produced and can contribute to regulatory learning: (1) self-assessment by a provider or deployer, (2) assessment within an AIRS, and (3) conformity assessment performed by a notified body.
@@
Figure 2. The flow of information in the self-assessment by an SME.
The flow of this scenario, with respect to the regulatory learning process, is shown in Figure 2. The legal requirements represent the macro-level pressure exerted on the micro-level SME to guarantee compliance. For the SME, legal requirements such as fairness are abstract with respect to their AI system. Hence, this step in the process requires careful interpretation by the SME in order to translate the legal requirements into appropriate technical requirements and a subsequent selection of corresponding assessment solutions. As provided for by the AI Act, said translation should be facilitated by regulatory instruments such as the harmonised standards that enable presumption with conformity.
The current absence of such regulatory instruments gives this selection particular potential for impact on regulatory learning. At the micro level, the iterative assessment of the AI system and the accompanying selection of tools with a consistent, reproducible methodology in an AITS makes the development of the AI system evident and transparent, potentially helping with interpretation of the legal requirements and the assessment results. At the same time, the scenario is not necessarily constrained to the micro level and the data generated may be passed onto the meso level. In fact, the SME may be a member of the Advisory Forum as provided for by Article 67(2), thus creating a direct feedback channel at the meso level to integrate the learned micro-level evidence. As the Advisory Forum collaborates with the AI Board, AI Office, and, if requested, with the Commission, the generated evidence can also reach the macro level. In addition, the SME may also participate in the standardisation processes of CEN and CENELEC (66), allowing them to directly affect the design of European standards, which constitutes another feedback loop within the regulatory learning process. In these processes, the assessment tool, method, and metrics selection helps the standardisation bodies detect potential candidates for adoption in standards.
AI Regulatory Sandbox. Some providers, especially startups and SMEs, do not have the resources or the legal expertise to perform the self-assessment and thus need external support. For those cases, and as a formal mechanism of regulatory learning, article 57 of the AI Act defines the previously mentioned AIRSes.
@@
In contrast to the self-assessment, however, this scenario represents not an indirect, but a direct, legally formal flow of information from the micro to the meso level. If the AITS methodology is employed on a consistent basis in AIRS engagements by the MSA, engagements become comparable and the MSA themselves gather evidence to improve their understanding of the pipeline from high-level legislation to technical operationalisation. As the MSA are tasked with sharing the collected data in their collaboration with the AI Board and the AI Office, it proliferates through the meso level, ultimately reaching the macro level. With the number of AIRS engagements growing, the machine-readable data generated within AITSes supports aggregation and allows scalable analysis at the meso and macro level. For the AI Office, it becomes particularly valuable in the design of guidelines and Codes of Practice, whereas the Commission may use insights to judge the appropriateness of standards to receive legal force through delegated acts.
@@
Conformity assessment by a notified body. The third, and last, scenario that we consider here consists in the certification process performed by a NB, certified by a national NA. This scenario, anchored in the AI Act via articles 43(1)(b), 43(3), and annex VII, can be considered a continuation of the two previous scenarios. Since the certification process consists uniquely in a review of the quality management system implemented and the technical documentation drafted by a provider or deployer, either self-assessment or an engagement similar to an AIRS has to precede it. Nevertheless, due to their responsibilities in reporting to the MSA, the NB serve as a first meso-level aggregator passing on the collected evidence. As a first information processor, they can perform a preliminary analysis of sector-specific data, improving clarity in the later stages of the regulatory learning process. For that to happen efficiently and at scale, manual, text-based compliance reports need to be replaced by specifically designed, standardised machine-readable formats, employed by AITSes.
4. Building Effective AITSes for Regulatory Learning
@@
The previous section examined the role of AITSes as a source of micro-level evidence in support of regulatory learning across different AI assessment scenarios. That discussion remained at an abstract, functional level. This section therefore focuses on the infrastructural and formal requirements needed to operationalise AITSes as a foundational component of the EU AI regulatory landscape. In light of the absence of mandated machine-readable formats in the AI Act (Schrepel, 2025), particular attention is given to solutions that enable scalable and (semi-)automated processing, with the aim of reducing the regulatory burden associated with manual documentation. While not exhaustive, the components discussed below are considered essential to achieving the flexibility required for this purpose.
@@
Extensible Formal Configuration Language. In (Buscemi et al., 2025), a Domain-Specific Language (DSL) was proposed as a formal, machine-readable syntactical specification framework for AITSes. From a legal perspective, this DSL can be understood as a standardised formal template that structures how technical requirements, assessment procedures, and evidential outputs are specified and documented. Acting as the common specification framework for AITSes, the DSL encodes assessment logic in a way that mirrors how AI systems are designed and evaluated in practice. It distinguishes between a stable core layer, which captures requirements that apply horizontally across sectors, and sector-specific extensions that introduce additional obligations relevant to particular application domains. This layered structure follows a piggybacking design pattern (Mernik et al., 2005), whereby sector-specific provisions are formally anchored to a shared core. As a result, domain-specific requirements remain legally and technically traceable to the horizontal foundations defined in the core DSL. This traceability supports regulatory learning both within individual sectors and across the regulatory system as a whole.
@@
Internal Unified Data Model. A core requirement of modular AITSes is the ability to integrate assessment solutions from an open catalogue, akin to a service registry (Papazoglou, 2003), and to present their results through harmonised reporting and visualisation mechanisms. These elements are essential for conformity assessment and regulatory learning, as they enable auditability, comparability, and ex post review of the technical evidence produced. In practice, assessment tools reflect the diversity of AI systems covered by the AI Act, ranging from traditional machine learning systems to generative models (European Commission, 2025a). Differences in system architectures, inputs, and outputs lead to heterogeneous evaluation methods and metrics, which complicates the harmonised storage and interpretation of results within an AITS. To preserve evidentiary value, the underlying data infrastructure must therefore support flexible data models and fine-grained access, including instance-level evidence where aggregate metrics would be insufficient (Burnell et al., 2023). Because assessment tools implement distinct evaluation logics, their integration requires not only architectural flexibility but also an explicit mapping between tool-specific outputs and a common evidentiary structure. To support a scalable ecosystem of assessment solutions while preserving human oversight, this mapping should be supported through semi-automated, user-validated mechanisms, providing an essential boundary resource enabling a plug-in architecture consistent with human-in-the-loop requirements (Ghazawneh and Henfridsson, 2013).
@@
Assessment tool documentation. Existing surveys and taxonomies of AI trustworthiness assessment reveal a highly fragmented evaluation landscape (Huang et al., 2025; Campagner et al., 2023; Newman, ; Braga et al., 2025). This fragmentation raises not only technical concerns but also legal ones, as divergent methodologies, metrics and documentation practices complicate the interpretation, comparability and auditability of assessment outcomes. Assessment tools may address the same legal requirement while relying on different underlying metrics or assumptions, thereby undermining consistent compliance assessment and legal certainty. The mere integration of such tools as plug-ins, while essential for flexibility, is therefore insufficient to meet the requirements of transparency and traceability. A complementary approach lies in the adoption of standardised reporting formats based on a shared vocabulary for assessment solutions. Comparable instruments already exist in the form of model cards and data sheets (Mitchell et al., 2019; Gebru et al., 2021; Roman et al., 2024), with more recent proposals such as AI cards, use case cards and AI product cards responding to documentation needs arising under the AI Act (Golpayegani et al., 2024; Hupont et al., 2024; Celepija et al., 2025).
@@
When accompanied by a machine-readable representation, such standardised formats enable the structured cataloguing of assessment solutions and support their consistent application across regulatory contexts. This, in turn, facilitates automated or semi-automated tool selection aligned with applicable legal requirements, reducing barriers to effective conformity assessment. At the meso and macro levels, standardised reporting also provides a basis for regulatory learning. By enabling systematic analysis of how assessment tools are applied in practice, competent authorities, the AI Office and standardisation bodies can derive evidence to inform guidance, standards and recommended practices, thereby strengthening consistency and legal certainty over time.
Reference Ontology of Metrics. While the aforementioned DSL provides a machine-readable framework for the syntactic structure of AITS specifications, it must be complemented by a reference ontology to resolve the real-world semantics of the domain (de Carvalho et al., 2014). Without this semantic layer, the DSL remains susceptible to interpretation errors across different sandbox instances; for example, if the same metric name is instantiated with divergent mathematical definitions, the resulting evidence becomes incomparable. To prevent such semantic heterogeneity, metric definitions, mapped to specific legal requirements and system types, should be formalised through machine-readable ontologies in open standards such as Resource Description Framework (RDF) (61). This hybrid approach mirrors established patterns in software engineering, where combining ontologies with DSLs has been proven to effectively manage the complexity of domains such as network configuration and validation (Miksa et al., 2010).
@@
In literature, significant steps in this direction have been made through ontologies such as AIRO, an ontology of risks in AI systems (Golpayegani et al., 2022) and semantic mapping between trustworthiness requirements and standards (Hernandez et al., 2025). However, the proposed ontologies remain high-level, adhering to the abstract definitions of the AI Act, with a closer link to operational technical requirements still missing. In (Di Scala et al., 2025), a conceptual language bridging the normative legal requirements to the functional components of an AI system is proposed. A combination of such semantic frameworks with standardised assessment tool documentation could drive the semi-automated instantiation of AITSes and the creation of auditable and machine-readable assessment reports with the needed consistency for analysis and regulatory learning at scale.
@@ 5. Discussion
The theoretical framework described in Section 3.1 establishes the micro level as the location of evidence generation, with the subsequent sections elaborating on AITSes as the critical technical infrastructure to complement formal regulatory mechanisms such as AIRSes. In this section, we discuss the implications that AITSes as standardised assessment tools would entail for regulatory learning, namely its strengths and associated socio-technical and socio-political challenges.
@@
First, a fundamental issue lies in learning at each level taking place at different speeds, with the learning at the micro level following the fast cycles of software development and thus vastly outpacing the learning at the macro level, itself constrained by the defined legislative review periods of up to four years (Article 112) (Ahern, 2021). The meso level, tasked with bridging this gap through harmonisation, currently risks becoming the bottleneck due to additional delays in the standardisation process. In this context, AITSes act as a technical buffer decoupling innovation speed from regulatory inertia. While the AI Act prescribes manual documentation, impeding scalable analysis, the proposed machine-readable formats of the AITSes provide a basis for future automated evidence aggregation. Due to data normalisation at the source, i.e. the micro level, through its internal unified data model, the use of AITSes streamlines the aggregation and interpretation of reported results at the meso level, alleviating the burden of interpreting manually reported documents at scale. This would in turn enable the meso level to identify technical trends with an agility that the traditional legislative cycles cannot match.
@@
A further, fundamental, challenge lies in the definition of adequate metrics and assessment methodologies to accurately capture the impact of an AI system on the fundamental rights the AI Act aims to codify. Abstract requirements such as fairness are inherently difficult to define at a technical level. This puts particular demands on the regulatory learning cycle as technical developments have to be taken into account when operationalising legislation. AITSes, through the envisioned plug-in architecture, can alleviate some of this pressure by enabling consistent experimentation and validation of diverse metrics. This way, empirical discovery of potentially adequate metrics, methods, and methodologies becomes available at a large scale, not just in the conformity assessment, but in the learning processes. Especially the meso actors, tasked with the translation of the legal requirements, can benefit from said discovery in the design of the sector-specific, vertical regulatory instruments such as Codes of Practice and the harmonised standards. At the same time, this particular concern calls for caution, as scholars have already warned against standards aiming at becoming too granular through the definition of, e.g., overly specific thresholds (Gornet and Maxwell, 2024).
@@
Beyond this normative challenge, the ambition of an adaptive governance model is not merely a technocratic matter in which learning naturally takes place from data. Instead, regulatory learning is inherently political, subject to significant influence from the interests of the involved stakeholders and decision makers. At the legislative level, this manifests as a question of political willingness. This issue is further compounded at the meso level where the data are aggregated and filtered before passing it to the macro level, especially in the standardisation efforts intended to provide vertical governance. As noted by Metcalf et al. (Metcalf, 2025), standardisation processes often involve stakeholders with asymmetric resources and capacities, which creates a non-negligible risk of regulatory capture. The researchers identify that, in such contexts, the interpretation of legal requirements into technical operationalisations may tend to prioritise feasibility, scalability, or market compatibility, potentially at the expense of more robust protections of fundamental rights. In particular the interests of the micro-level actors, such as SMEs, should be given more weight (Kilian et al., 2025), a step that the AI Act has already taken through their inclusion in the regulatory learning process via the Advisory Forum. As AITSes represent a technical infrastructure enabling transparent, reproducible, and consistent assessments, it is decidedly not prescriptive. It is a tool enabling generation of data, with political issues out of its scope. However, the AITS methodology and subsequent semi-automatic reporting would alleviate the need for resources and effectively democratise the aggregation layer through potentially higher impact by smaller stakeholders. Requiring the use of the evidential data in the standardisation process could potentially transform the decision making by giving equal weights to the input from small actors as the one from large industry players.
@@
Finally, evidence generation by providers and deployers is insufficient as the data need to be shared to enable learning at the higher levels. This raises the crucial question concerning the incentive for industry stakeholders to allow access to their logged data. On the one hand, privacy and data protection concerns naturally arise as companies will not want to reveal internal information. On the other hand, there is a risk that data generated within an AITS could be perceived as self-incriminating, discouraging its use as a tool for compliance support rather than post-hoc accountability. To mitigate such liability concerns, early stage AITS findings could, similarly to the AIRS engagements, be legally treated as preliminary learning signals instead of grounds for potential sanctions, acknowledging the iterative nature of development and assessment. This could lead to actors at the micro level being more open towards sharing their results arising from self-assessment beyond just going through the necessary steps for conformity assessment, as it would signal that their input is explicitly desired to review the legislation and regulations. Furthermore, by reducing the resources required in drafting evidence of compliance, AITSes could see improved adoption in the industry, fuelling both enforcement and learning at the same time.